OpenStudy (kkrpbra):

can someone show 3 examples of equations

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 2x - 3 = 6x - 9 }\) \(\Large \color{MidnightBlue}{\Rightarrow x + 4 = y * 8 }\) \(\Large \color{MidnightBlue}{\Rightarrow a + b + c = 2 + 6 + a }\) Though, not all of them can be solved, but these are equations.